### v1

# Unified model for non-Abelian braiding of Majorana and Dirac fermion zero modes

Abstract
Majorana zero modes (MZMs) are the most intensively studied non-Abelian anyons. The Dirac fermion zero modes in topological insulators can be interpreted as the symmetry-protected “doubling” of the MZMs, suggesting an intrinsic connection between the quantum statistics of the two types of zero modes. Here we find that the minimal Kitaev chain model provides a unified characterization of the non-Abelian braiding statistics of both the MZMs and Dirac fermion zero modes under different parameter regimes. In particular, we introduce a minimal tri-junction setting based on the minimal Kitaev chain model and show it facilitates the unified scheme of braiding Dirac fermion zero modes, as well as the MZMs in the assistance of a Dirac mode. This unified minimal model unveils that the non-Abelian braiding of the MZMs can be continuously extended to the realm of Dirac fermion zero modes. The present study reveals deeper insights into the non-Abelian statistics and enables a broader search for non-Abelian anyons beyond the scope of MZMs.
preprint: APS/123-QED
Introduction. Anyons Wilczek (1982) obeying fractional statistics are exotic quasiparticles living in (2+1)-dimensional space. In condensed matter physics, anyons usually emerge as composites formed from charged particles (like electrons) and vortices. This anyon model was raised by Wilczek in 1982 Wilczek (1982) and first exemplified by the Laughlin quasiparticles Laughlin (1983) in the fractional quantum Hall (FQH) system Tsui et al. (1982). Also in the $\nu=5/2$ FQH state Willett et al. (1987), the more intriguing non-Abelian anyons are proposed to emerge as the elementary excitations binding with half-vortices Moore and Read (1991); Wen (1991). This type of excitation in correlated electron systems also manifests itself as a half-vortex-bound state in $p$-wave superconductors Ivanov (2001), which is well known as the Majorana zero mode (MZM) Kitaev (2001).
Following these pioneering works, great efforts have been devoted to seeking the experimental evidence of MZMs because of their great potential for fault-tolerant topological quantum computation Kitaev (2003); Nayak et al. (2008). MZMs have been proposed to appear in a number of platforms, an incomplete list includes superconductor-semiconductor heterostructures Fu and Kane (2008); Mourik et al. (2012); Das et al. (2012); Cao et al. (2023), ferromagnetic atomic chains on a superconductor Nadj-Perge et al. (2014), and vortices in topological superconductors Sun et al. (2016); Wang et al. (2018); Zhang et al. (2018); Liu and Ding (2023). However, the complexity of the superconductor-proximated heterostructure Ji and Wen (2018); Cao et al. (2023) as well as the possible trivial explanations for the transport signals Liu et al. (2012, 2017) still hinder the appearance of the smoking-gun experimental signals. An altenative way is to seek the symmetry protected Majorana pairs. The most studied one among them is the Majorana Kramer’s pair (MKP) Wong and Law (2012); Zhang et al. (2013), where a pair of MZMs is related by time-reversal (TR) symmetry. Owing to the TR symmetry protection, MKPs have been shown to obey non-Abelian braiding, rendering a novel type of symmetry-protected non-Abelian statistics Liu et al. (2014); Wölms et al. (2014, 2016); Gao et al. (2016), as well as topological quantum computing Schrade and Fu (2022).
When the TR symmetry (an anti-unitary symmetry) in the MKP is replaced by a unitary symmetry, the symmetry-protected non-Abelian statistics remains Hong et al. (2022). This kind of Majorana pair related by a unitary symmetry is equivalent to a zero-energy Dirac fermion mode bound to a half vortex Wu et al. (2020a); Yasui et al. (2012), and the proof of the non-Abelian statistics for Dirac fermion modes can be traced back to Yasui et al.’s work in 2012 Yasui et al. (2012). This kind of zero-energy Dirac fermion mode can appear as a bound state with a half-vortex attached in a two-dimensional topological insulator Shan et al. (2011). The incarnation of this non-Abelian Dirac fermion zero mode in one-dimensional system is the celebrated Jackiw-Rebbi zero mode, whose half-integer fractional charge Jackiw and Rebbi (1976); Su et al. (1979) and non-Abelian braiding Klinovaja and Loss (2013, 2015); Boross et al. (2019) have been discussed in the early works.
The fact that the Dirac fermion zero modes obeying non-Abelian braiding statistics can be understood in the following way. Given the presence of a unitary symmetry, a set of Dirac fermion zero modes can be decomposed into two sets of MZMs in which the cross coupling between the opposite sets is prohibited by symmetry. Therefore, the braiding of Dirac fermion modes can be characterized as the direct product of the braiding of two sets of MZMs Hong et al. (2022). In this way, the non-Abelian braiding statistics of the Dirac fermion zero modes are essentially related to the pairs of MZMs. In addition, it has long been discussed that the braiding operation of MZMs can be facilitated by a Dirac fermion zero mode, which is known as the “quantum-dot (QD)-assisted Majorana braiding” Flensberg (2011); Liu et al. (2021); Gong et al. (2022); Krøjer et al. (2022); Xu et al. (2023). In this braiding scheme, the low-lying states involved include both the MZMs and Dirac fermion modes. The presence of such braiding scheme intermediating between the Majorana and the Dirac fermionic systems further hints us that these braiding schemes mentioned above might be described in a unified model.
In this Letter, we demonstrate that the minimal Kitaev chain model Leijnse and Flensberg (2012); Tsintzis et al. (2024) can play this role. To be specific, under different parameter choices, the minimal Kitaev chain can depict the non-Abelian braiding of MZMs, Dirac fermion zero modes, as well as the QD-assisted Majorana braiding. Significantly, the quasiparticles being braided in our setting can switch from the MZMs to Dirac fermion zero modes with a small change in parameters. It indicates that the search for non-Abelian anyons shall not be confined to the realm of MZMs. In fact, the degenerate Dirac fermion zero modes can always be regarded as the symmetry protected “doubling” of the MZMs Hong et al. (2022, 2024). In addition to providing new insights into non-Abelian anyons, our proposed study based on minimal Kitaev chain model is experimentally feasible and could be detected using ultracold atoms or QD arrays.
[…]
(a) Refer to caption (b) Refer to caption
Figure 1: (a) Sketch of the Y-junction jointed by minimal Kitaev chains. (b) The braiding operation corresponds to a closed loop (started and ended at the point marked in red) in the parameter space.
However, properly defining non-Abelian braiding requests at least four MZMs Ivanov (2001). Therefore, in order to perform the braiding of the PMMs, the minimal device required is a four-site Y-shaped junction Alicea et al. (2011); Sau et al. (2011); von Oppen et al. (2017) jointed by three minimal Kitaev chains [see Fig. 1(a)]
$H=\sum\limits_{j=1}^{3}\chi_{j}d_{0}^{\dagger}d_{j}-\chi_{j}^{*}d_{0}d_{j}^{\dagger}+\Omega_{j}d_{0}^{\dagger}d_{j}^{\dagger}-\Omega_{j}^{*}d_{0}d_{j}.$ (1)
Here, $d_{j}$ ($d_{j}^{\dagger}$) annihilates (creates) an electron on site $j$, while $d_{0}$ ($d_{0}^{\dagger}$) annihilates (creates) an electron on the central site, $\chi_{j}$ and $\Omega_{j}$ are the amplitudes of the hopping and superconducting pairing, respectively. In general, $\chi$’s and $\Omega$’s are complex as $\chi_{j}=|\chi_{j}|e^{i\theta_{\chi_{j}}}$ and $\Omega_{j}=|\Omega_{j}|e^{i\theta_{\Omega_{j}}}$. Significantly, it is impossible to gauge away all these six phases $\theta_{\chi_{j}}$ and $\theta_{\Omega_{j}}$ ($j=1,2,3$), since in Eq. (1), there are only four electron operators can be utilized to absorb the gauge phases. In other words, the “sweet spot” condition cannot always be satisfied when we consider the braiding in this minimal-Kitaev-chain-based Y-junction. Nonetheless, the non-Abelian braiding physics shall not collapse due to a slight deviation of the parameters from the “sweet spot”. Meanwhile, the low-lying states of this Y-junction being braided now are not necessarily PMMs. It implies that this Y-junction can depict the non-Abelian braiding of different kinds of low-lying modes for different choices of $\chi$’s and $\Omega$’s. Specifically, there are four distinct cases based on the magnitudes and the phases of $\chi_{j}$ and $\Omega_{j}$, which are summarized in TABLE 1 and will be discussed below in detail.
$\theta_{\chi_{j}},\theta_{\Omega_{j}}=0$ $\theta_{\chi_{j}},\theta_{\Omega_{j}}\neq 0$
$|\chi_{j}|=|\Omega_{j}|$ Case I: poorman’s MZMs (PMMs) Case II: QD-assisted MZM braiding
$|\chi_{j}|\neq|\Omega_{j}|$ Case III: Dirac fermion zero modes (unitary-symmetry preserved) Case IV: Dirac fermion zero modes (unitary-symmetry broken)
Table 1: The minimal-Kitaev-chain-based Y-junction can depict the non-Abelian braiding of different types of low-lying modes depending on the choices of $\chi$’s and $\Omega$’s.
The braiding operation in this Y-junction is conducted through coupling the three outer sites with the central site in turn Sau et al. (2011); von Oppen et al. (2017). Without loss of generality, we choose swapping the low-lying modes localized at site 2 and site 3 of this Y-junction [see Fig. 1(a)] as an instance. The Y-junction is initialized by setting $\chi_{2}$, $\chi_{3}$, $\Omega_{2}$, and $\Omega_{3}$ to zero, while setting $\chi_{1}$ and $\Omega_{1}$ to finite values. In this way, both site 2 and site 3 become isolated so that each of them contains a zero-energy mode. In the first step of the braiding, $\chi_{2}$ and $\Omega_{2}$ are adiabatically increased to finite values, while $\chi_{1}$ and $\Omega_{1}$ are adiabatically decreased to zero. Here, we assume the ratios $|\chi_{1}|/|\Omega_{1}|$ and $|\chi_{2}|/|\Omega_{2}|$ remain constants during the braiding, and the phases $\theta_{\chi_{j}},\theta_{\Omega_{j}}$ also remain unchanged. In this way, the zero-energy mode initially localized at site 2 is adiabatically moved to site 1. In the second step, in the same manner, $\chi_{3}$ and $\Omega_{3}$ are both increased to finite values, and $\chi_{2}$ and $\Omega_{2}$ are decreased to zero in an adiabatic way so that the zero-energy mode in site 3 is moved to site 2. In the final step, the braiding is accomplished by increasing $\chi_{1}$ and $\Omega_{1}$ to their initial values while decreasing $\chi_{3}$ and $\Omega_{3}$ to zero. Such a braiding process corresponds to a closed loop in the parameter space [see Fig. 1(b)].
Non-Abelian braiding of MZMs. We firstly focus on the braiding under the conditions $|\chi_{j}|=|\Omega_{j}|$ ($j=1,2,3$), i.e. case I and case II in TABLE 1. By decomposing the electron operators $d_{j}$ into Majorana operators $\gamma_{j}$ and $\widetilde{\gamma}_{j}$ ($j=0,1,2,3$), the original Hamiltonian [Eq. (1)] can be rewritten as Sup :
$H$ $=$ $-i|\chi_{1}|\gamma_{0}\gamma_{1}-i|\chi_{2}|(-\sin\varphi_{2}\widetilde{\gamma}_{0}+\cos\varphi_{2}\gamma_{0})\gamma_{2}$ (2)
$-i|\chi_{3}|(-\sin\varphi_{3}\widetilde{\gamma}_{0}+\cos\varphi_{3}\gamma_{0})\gamma_{3}.$
Here, as we have stated before, $\varphi_{2}\equiv(\theta_{\chi_{2}}+\theta_{\Omega_{2}}-\theta_{\chi_{1}}-\theta_{\Omega_{1}})/2$ and $\varphi_{3}\equiv(\theta_{\chi_{3}}+\theta_{\Omega_{3}}-\theta_{\chi_{1}}-\theta_{\Omega_{1}})/2$ are two phase constants ($\theta_{\chi_{j}}$ and $\theta_{\Omega_{j}}$ remain unchanged during the braiding) that cannot be absorbed even after gauge transformation Sup . This Majorana Hamiltonian [Eq. (2)] can be schematically presented by Fig. 2(a). Specifically, we consider the parity-even subspace spanned by six MZMs $\gamma_{0}$, $\gamma_{1}$, $\gamma_{2}$, $\gamma_{3}$, $\widetilde{\gamma}_{0}$, and $\widetilde{\gamma}_{1}$ [dashed box in Fig. 2(a)], because fermion parity is conserved here and the Hamiltonian in the parity-even subspace is identical to that in the parity-odd subspace (up to a constant) Sup . It is worth noting that though $\widetilde{\gamma}_{1}$ is isolated from the other MZMs, it is still intentionally considered to ensure that the dimension of the Fock space is an integer.
In the case I of TABLE 1 that $\varphi_{2},\varphi_{3}=0$, the two-fold degenerate negative-energy states are gapped from the two-fold degenerate positive-energy states (the gap is denoted by $\Delta$), and these degeneracies are preserved throughout the braiding [see Fig. 2(b)]. In the adiabatic limit, where the braiding time $T\gg 1/\Delta$, it is meaningful to consider only the two-fold degenerate ground state manifold. The Berry phase during the braiding can be evaluated analytically Sup , which leads to an evolution operator as $U_{\mathrm{case~{}I}}=\begin{pmatrix}e^{-i\pi/4}&0\\ 0&e^{i\pi/4}\end{pmatrix}$. This is precisely the Majorana braiding operator Ivanov (2001); von Oppen et al. (2017) as expected, since the “sweet spot” condition is satisfied in this case, ensuring that all the low-lying modes in the system are PMMs.
Figure 2: (a) Schematic representation of the Majorana Hamiltonian [Eq. (2)] under the conditions that $|\chi_{j}|=|\Omega_{j}|$. The Fock space we consider is spanned by six MZMs inside the dashed box. (b), (c) The energy spectra during the braiding as (b) Case I in TABLE 1 that $|\chi_{j}|=|\Omega_{j}|$ and $\varphi_{2},\varphi_{3}=0$. All the states in (b) are doubly degenerate; (c) Case II in TABLE 1 that $|\chi_{j}|=|\Omega_{j}|$ and $\varphi_{2},\varphi_{3}\neq 0$ ($\varphi_{2}=\pi/6$, and $\varphi_{3}=\pi/12$ here for illustration). $|\chi|\equiv\sqrt{|\chi_{1}|^{2}+|\chi_{2}|^{2}+|\chi_{3}|^{2}}$ is chosen as the energy unit.
In the case II of TABLE 1 that $\varphi_{2},\varphi_{3}\neq 0$, the two-fold ground-state degeneracy is lifted except for three special parameter points during the braiding [see Fig. 2(c)]. Therefore, in addition to the Berry phase, the dynamic phase part in the evolution operator $U=\hat{P}\exp[\int-iE(t)\mathrm{d}t+i\mathbf{A}(\boldsymbol{\lambda})\cdot\mathrm{d}\boldsymbol{\lambda}]$ Wilczek and Zee (1984); Chang (2021) generally cannot be dropped. Here, $\hat{P}$ represents the time-ordering operator, $E$ is a diagonal matrix corresponding to the eigenenergy spectra, and $\mathbf{A}$ is the Berry connection. Under the adiabatic condition, where $T\gg 1/\Delta$, we could still focus only on the two lowest states with negative energies [see Fig. 2(c)]. The corresponding evolution operator can be evaluated numerically as $U_{\mathrm{case~{}II}}=e^{i\theta}\begin{pmatrix}e^{-i(\pi/4+\delta)}&0\\ 0&e^{i(\pi/4+\delta)}\end{pmatrix}$. Here, the average energy of these two lowest states contributes a trivial overall dynamic phase $\theta$, while the energy difference between these two lowest states leads to a dynamic phase difference $\delta$ Sup . If $\frac{1}{\varphi_{2}},\frac{1}{\varphi_{3}}\gg T\gg\frac{1}{\Delta}$, then this dynamic phase difference $\delta$ is relatively small so now the braiding operator comes back to the form of $\mathrm{diag}\{e^{-i\pi/4},e^{i\pi/4}\}$. It indicates that the geometric phase accumulated remains $\mp\pi/4$ even when $\varphi_{2},\varphi_{3}\neq 0$, which can be confirmed by analytically evaluating the geometric phase part of the evolution operator $\hat{P}\exp[\int i\mathbf{A}(\boldsymbol{\lambda})\cdot\mathrm{d}\boldsymbol{\lambda}]$ Sup . Otherwise, the dynamic phase dominates when $\frac{1}{\varphi_{2}},\frac{1}{\varphi_{3}}\ll T$, causing the braiding results to oscillate with increasing $T$ Sup . Such results that in addition to the geometric phase of $\mp\pi/4$, an undesired dynamic phase being also involved has been observed in the QD-assisted braiding Flensberg (2011); Liu et al. (2021); Gong et al. (2022); Xu et al. (2023). In fact, as shown in Fig. 2(a), two MZMs $\gamma_{2}$ and $\gamma_{3}$ being braided here are swapped via coupling to both the two Majorana components $\gamma_{0}$ and $\widetilde{\gamma}_{0}$ located at the central site. Therefore, the minimal Kitaev chain here exactly describes the QD-assisted braiding that two MZMs are swapped via a Dirac fermion zero mode.
Non-Abelian braiding of Dirac fermion zero modes. The more intriguing scenario is that even though the equalities $|\chi_{j}|=|\Omega_{j}|$ ($j=1,2,3$) are only slightly violated, the low-lying modes being braided are no longer single MZMs, but Majorana pairs. To be specific, under the conditions $|\chi_{j}|\neq|\Omega_{j}|$ ($j=1,2,3$), i.e. cases III and IV in TABLE 1, the original Hamiltonian [Eq. (1)] in the Majorana form reads Sup
$H$ $=$ $-i\frac{|\chi_{1}|+|\Omega_{1}|}{2}\gamma_{0}\gamma_{1}-i\frac{|\chi_{1}|-|\Omega_{1}|}{2}\widetilde{\gamma}_{0}\widetilde{\gamma}_{1}$ (3)
-- $i\cos\varphi_{2}\bigr{(}\frac{|\chi_{2}|+|\Omega_{2}|}{2}\gamma_{0}\gamma_{2}+\frac{|\chi_{2}|-|\Omega_{2}|}{2}\widetilde{\gamma}_{0}\widetilde{\gamma}_{2}\bigr{)}$
-- $i\cos\varphi_{3}\bigr{(}\frac{|\chi_{3}|+|\Omega_{3}|}{2}\gamma_{0}\gamma_{3}+\frac{|\chi_{3}|-|\Omega_{3}|}{2}\widetilde{\gamma}_{0}\widetilde{\gamma}_{3}\bigr{)}$
$+$ $i\sin\varphi_{2}\bigr{(}\frac{|\chi_{2}|+|\Omega_{2}|}{2}\widetilde{\gamma}_{0}\gamma_{2}-\frac{|\chi_{2}|-|\Omega_{2}|}{2}\gamma_{0}\widetilde{\gamma}_{2}\bigr{)}$
$+$ $i\sin\varphi_{3}\bigr{(}\frac{|\chi_{3}|+|\Omega_{3}|}{2}\widetilde{\gamma}_{0}\gamma_{3}-\frac{|\chi_{3}|-|\Omega_{3}|}{2}\gamma_{0}\widetilde{\gamma}_{3}\bigr{)}.$
One can notice that for case III in TABLE 1 that $\varphi_{2}=\varphi_{3}=0$, the cross-coupling terms like $i\gamma_{i}\widetilde{\gamma}_{j}$ vanish. In this way, a unitary symmetry $\mathcal{R}$ defined as $\mathcal{R}_{ij}i\gamma_{i}\widetilde{\gamma}_{j}\mathcal{R}_{ij}^{-1}=-i\gamma_{i}\widetilde{\gamma}_{j}$ is preserved Hong et al. (2022), and all these MZMs are separated into two sets as shown in Fig. 3(a). According to the braiding protocol introduced above, these two independent sets of MZMs $\{\gamma_{i}\}$ and $\{\widetilde{\gamma}_{i}\}$ are braided simultaneously. It is worth noticing that $\gamma_{i}$ and $\widetilde{\gamma}_{i}$ are spatially fully-overlapped. Therefore, the low-lying modes being braided here are actually fully-overlapped Majorana pairs respecting local unitary symmetry $\mathcal{R}$. Then each Majorana pair as a whole is actually Dirac fermion zero mode Hong et al. (2022); Wu et al. (2020a, 2023) that the corresponding braiding properties can be constructed from the Majorana ones.
In the adiabatic limit $T\gg\frac{1}{w},\frac{1}{\Delta}$ [$w$ is the (average) energy difference between the lowest two-fold states and the second lowest two-fold states, see Fig. 3(c)], the Berry phase for the lowest two-fold states turns out to be the projection of the tensor product of the Majorana one’s into the parity-even subspace as ^$\hat{\mathbb{P}}_{\mathrm{even}}[\mathrm{diag}\{e^{-i\pi/4},e^{i\pi/4}\}\otimes\mathrm{diag}\{e^{-i\pi/4},e^{i\pi/4}\}]=\mathrm{diag}\{-i,i\}$. Similarly, the Berry phase for the second lowest two-fold states is ^$\hat{\mathbb{P}}_{\mathrm{even}}[\mathrm{diag}\{e^{-i\pi/4},e^{i\pi/4}\}\otimes\mathrm{diag}\{e^{i\pi/4},e^{-i\pi/4}\}]=\mathrm{diag}\{1,1\}$ Sup . In addition to this, the dynamic phases also contribute when $T\gg\frac{1}{w},\frac{1}{\Delta}$, and hence the braiding operator reads $U_{\mathrm{case~{}III}}=e^{i\theta}\left[e^{-i\delta}\begin{pmatrix}-i&0\\ 0&i\end{pmatrix}\oplus e^{i\delta}\begin{pmatrix}1&0\\ 0&1\end{pmatrix}\right]$. Here, $\theta$ ($\delta$) are the dynamic phases corresponding to the average energy of (energy difference between) the lowest and second lowest two-fold dengenerate states.
By contrast, for $1/w\gg T\gg 1/\Delta$, the dynamic phase of the four lowest states only contributes a trivial overall phase and can be dropped. Moreover, the Berry phase for the Majorana set $\{\widetilde{\gamma}_{j}\}$ now reduces to an identity $I_{4\times 4}$ since $\widetilde{\gamma}_{j}$’s are braided in the non-adiabatic limit $1/w\gg T$ Chang (2021). Therefore, in this case, the braiding operator is obtained by projecting the tensor product $\mathrm{diag}\{e^{-i\pi/4},e^{i\pi/4}\}\otimes I_{4\times 4}$ into the parity-even subspace, giving $U_{\mathrm{case~{}III}}=\mathrm{diag}\{e^{-i\pi/4},e^{-i\pi/4},e^{i\pi/4},e^{i\pi/4}\}$ Sup .
Figure 3: (a), (b) Schematic representation of the Majorana Hamiltonian [Eq. (3)] under the conditions of (a) Case III in TABLE 1; and (b) Case IV in TABLE 1. (c), (d) The energy spectra during the braiding as (c) Case III that $|\chi_{j}|\neq|\Omega_{j}|$ and $\varphi_{2},\varphi_{3}=0$ ($|\Omega_{i}|/|\chi_{i}|=0.8$ here for illustration); and (d) Case IV that $|\chi_{j}|\neq|\Omega_{j}|$ and $\varphi_{2},\varphi_{3}\neq 0$ ($|\Omega_{i}|/|\chi_{i}|=0.8$, $\varphi_{2}=\pi/6$, and $\varphi_{3}=\pi/12$ here for illustration). The energy unit $|\chi|\equiv\sqrt{|\chi_{1}|^{2}+|\chi_{2}|^{2}+|\chi_{3}|^{2}}$, and all the states in (c) and (d) are doubly degenerate.
Finally, for case IV in TABLE 1 that $\varphi_{2},\varphi_{3}\neq 0$, the unitary symmetry $\mathcal{R}$ is broken and the two sets of MZMs $\{\gamma_{i}\}$ and $\{\widetilde{\gamma}_{i}\}$ are coupled via the cross-coupling terms [see Fig. 3(b)]. The energy spectra in Fig. 3(d) shows that all the two-fold degeneracies are still preserved, while the energies of these degenerate states vary during the braiding. Through numerical calculation Sup , we have confirmed that the Berry phase accumulated in case IV is exactly the same as the one in case III (up to a gauge transform). This is in analogy with the consequence stated above that the Berry phase is the same for case II and case I. The difference lies in that now the dynamic phases $\theta\mp\delta$ are $\varphi_{2}$-, $\varphi_{3}$-, and $|\Omega_{j}|/|\chi_{j}|$-dependent since the energy during the braiding vary with $\varphi_{2}$, $\varphi_{3}$, and $|\Omega_{j}|/|\chi_{j}|$. Meanwhile, the braiding results still exhibit an oscillation with increasing $T$ Sup . Moreover, the (average) energy difference $w$ between the two-fold ground states and the second-lowest two-fold states also increases with $\varphi_{2}$, $\varphi_{3}$, and $\big{|}|\chi_{j}|-|\Omega_{j}|\big{|}$. In this way, the $1/w\gg T\gg 1/\Delta$ condition can no longer be satisfied when $\varphi_{2}$, $\varphi_{3}$, or $\big{|}|\chi_{j}|-|\Omega_{j}|\big{|}$ is significantly non-zero. As a result, an additional dynamic phase is always accumulated alongside the geometric phase, providing a clear signature of the symmetry-breaking effect Wu et al. (2020a, 2023, b) on the braiding of Dirac fermion zero modes.
Discussions. The minimal Kitaev chain model is experimentally feasible, in ultracold atom systems for instance, especially for $|\Omega_{j}|=0\neq|\chi_{j}|$, as the complexity of pairing term is circumvented. In this case, the low-lying modes are Dirac fermion modes with unitary symmetry preserved (case III) or broken (case IV), other than MZMs. In this condition, the Hamiltonian in Eq. (1) can be engineered by utilizing the internal degree of freedom of atoms. Taking 171Yb fermions as an example, the four braiding modes correspond to the atomic spin states, 1S0$|F=1/2,m_{F}=\pm 1/2\rangle$, and ${}^{3}P$0$|F=1/2,m_{F}=\pm 1/2\rangle$, which have significantly long lifetime. The tunnelings among the four fermion modes are highly controllable by manipulating Raman and clock transitions Jenkins et al. (2022).
Acknowledgements. This work was financially supported by the Innovation Program for Quantum Science and Technology (Grant No. 2021ZD0302400 and No. 2021ZD0302000), the National Natural Science Foundation of China (Grant No. 12304194, No. 12425401 and No. 12261160368), National Key Research and Development Program of China (2021YFA1400900), and by Shanghai Municipal Science and Technology (Grant No. 24DP2600100).

### v2

@@
 Majorana zero modes (MZMs) are the most intensively studied non-Abelian anyons.
-The Dirac fermion zero modes in topological insulators can be interpreted as the symmetry-protected “doubling” of the MZMs, suggesting an intrinsic connection between the quantum statistics of the two types of zero modes.
-Here we find that the minimal Kitaev chain model provides a unified characterization of the non-Abelian braiding statistics of both the MZMs and Dirac fermion zero modes under different parameter regimes.
+The Dirac fermion zero modes in topological insulators, which are symmetry-protected doubling of MZMs under fermion number conservation, offer an alternative approach to explore non-Abelian anyons.
+However, a unified model that elucidates the braiding statistics of these types of topological zero modes remains absent.
+We show that the minimal Kitaev chain model beyond fine-tuning regime provides a unified characterization of the non-Abelian statistics of both MZMs and Dirac fermion zero modes in different parameter regimes.
 In particular, we introduce a minimal tri-junction setting based on the minimal Kitaev chain model and show it facilitates the unified scheme of braiding Dirac fermion zero modes, as well as the MZMs in the assistance of a Dirac mode.
-This unified minimal model unveils that the non-Abelian braiding of the MZMs can be continuously extended to the realm of Dirac fermion zero modes.
-The present study reveals deeper insights into the non-Abelian statistics and enables a broader search for non-Abelian anyons beyond the scope of MZMs.
-preprint: APS/123-QED
+This unified minimal model provides deeper insights into non-Abelian statistics, demonstrating that the non-Abelian braiding of MZMs can be continuously extended to encompass Dirac fermion zero modes.
+The minimal Kitaev chain has been realized in coupled quantum dots [Nature 614, 445 (2023)].
+Our extension, which demonstrates novel nontrivial phases with non-Abelian MZM pairs and Dirac zero modes emerging in the broader parameter regimes without fine-tuning, expands the accessible experimental parameter space and enhances the feasibility of observing non-Abelian statistics in the minimal Kitaev chain model.
 Introduction.
-Anyons Wilczek (1982) obeying fractional statistics are exotic quasiparticles living in (2+1)-dimensional space.
-In condensed matter physics, anyons usually emerge as composites formed from charged particles (like electrons) and vortices.
-This anyon model was raised by Wilczek in 1982 Wilczek (1982) and first exemplified by the Laughlin quasiparticles Laughlin (1983) in the fractional quantum Hall (FQH) system Tsui et al.
+Anyons Wilczek (1982) in condensed matter systems usually emerge as composites formed from charged particles (like electrons) and vortices.
+This anyon model was raised by Wilczek Wilczek (1982) and first exemplified by the Laughlin quasiparticles Laughlin (1983) in the fractional quantum Hall (FQH) system Tsui et al.
 (1982).
@@
 (1987), the more intriguing non-Abelian anyons are proposed to emerge as the elementary excitations binding with half-vortices Moore and Read (1991); Wen (1991).
-This type of excitation in correlated electron systems also manifests itself as a half-vortex-bound state in $p$-wave superconductors Ivanov (2001), which is well known as the Majorana zero mode (MZM) Kitaev (2001).
-Following these pioneering works, great efforts have been devoted to seeking the experimental evidence of MZMs because of their great potential for fault-tolerant topological quantum computation Kitaev (2003); Nayak et al.
-(2008).
-MZMs have been proposed to appear in a number of platforms, an incomplete list includes superconductor-semiconductor heterostructures Fu and Kane (2008); Mourik et al.
+This type of excitation also manifests itself as a half-vortex-bound state in $p$-wave superconductors Ivanov (2001), which is well known as the Majorana zero mode (MZM) Kitaev (2001).
+Due to their significant potential for fault-tolerant topological quantum computation Kitaev (2003); Nayak et al.
+(2008), substantial efforts have been dedicated to seeking experimental evidence of MZMs in various platforms including superconductor-semiconductor heterostructures Fu and Kane (2008); Mourik et al.
 (2012); Das et al.
@@
 (2018); Liu and Ding (2023).
-However, the complexity of the superconductor-proximated heterostructure Ji and Wen (2018); Cao et al.
-(2023) as well as the possible trivial explanations for the transport signals Liu et al.
+However, the complexity of the superconductor-proximitized structure Ji and Wen (2018); Cao et al.
+(2023) and the possible trivial explanations for the transport signals Liu et al.
 (2012, 2017) still hinder the appearance of the smoking-gun experimental signals.
-An altenative way is to seek the symmetry protected Majorana pairs.
+An alternative approach for non-Abelian anyons is the symmetry-protected Majorana pairs.
 The most studied one among them is the Majorana Kramer’s pair (MKP) Wong and Law (2012); Zhang et al.
 (2013), where a pair of MZMs is related by time-reversal (TR) symmetry.
-Owing to the TR symmetry protection, MKPs have been shown to obey non-Abelian braiding, rendering a novel type of symmetry-protected non-Abelian statistics Liu et al.
+Owing to the TR symmetry protection, MKPs have been shown to obey a novel type of symmetry-protected non-Abelian statistics Liu et al.
 (2014); Wölms et al.
 (2014, 2016); Gao et al.
-(2016), as well as topological quantum computing Schrade and Fu (2022).
-When the TR symmetry (an anti-unitary symmetry) in the MKP is replaced by a unitary symmetry, the symmetry-protected non-Abelian statistics remains Hong et al.
-(2022).
-This kind of Majorana pair related by a unitary symmetry is equivalent to a zero-energy Dirac fermion mode bound to a half vortex Wu et al.
+(2016); Schrade and Fu (2022).
+When the TR symmetry (an anti-unitary symmetry) in the MKP is replaced by a unitary symmetry, this type of Majorana pair will undergo a continuous crossover to a zero-energy Dirac fermion mode bound to a half vortex Wu et al.
 (2020a); Yasui et al.
-(2012), and the proof of the non-Abelian statistics for Dirac fermion modes can be traced back to Yasui et al.’s work in 2012 Yasui et al.
+(2012); Hong et al.
+(2022) when fermion number conservation is imposed Hong et al.
+(2022, 2024).
+Here, the attached half vortex is regarded as the origin of the non-Abelian statistics of these Dirac fermion zero modes Yasui et al.
 (2012).
 This kind of zero-energy Dirac fermion mode can appear as a bound state with a half-vortex attached in a two-dimensional topological insulator Shan et al.
-(2011).
-The incarnation of this non-Abelian Dirac fermion zero mode in one-dimensional system is the celebrated Jackiw-Rebbi zero mode, whose half-integer fractional charge Jackiw and Rebbi (1976); Su et al.
-(1979) and non-Abelian braiding Klinovaja and Loss (2013, 2015); Boross et al.
-(2019) have been discussed in the early works.
-The fact that the Dirac fermion zero modes obeying non-Abelian braiding statistics can be understood in the following way.
-Given the presence of a unitary symmetry, a set of Dirac fermion zero modes can be decomposed into two sets of MZMs in which the cross coupling between the opposite sets is prohibited by symmetry.
-Therefore, the braiding of Dirac fermion modes can be characterized as the direct product of the braiding of two sets of MZMs Hong et al.
-(2022).
-In this way, the non-Abelian braiding statistics of the Dirac fermion zero modes are essentially related to the pairs of MZMs.
-In addition, it has long been discussed that the braiding operation of MZMs can be facilitated by a Dirac fermion zero mode, which is known as the “quantum-dot (QD)-assisted Majorana braiding” Flensberg (2011); Liu et al.
+(2011), or as the celebrated Jackiw-Rebbi zero mode in one-dimensional system, whose $1/2$ fractional charge Jackiw and Rebbi (1976); Su et al.
+(1979) and fractional statistics Klinovaja and Loss (2013, 2015); Boross et al.
+(2019) were discussed in early studies.
+Although the braiding statistics of the Dirac fermion zero modes has an intrinsic relation to MZM pairs Hong et al.
+(2022), a unified model that simultaneously realizes non-Abelian statistics of MZMs and Dirac fermion zero modes is lacking.
+Currently, the three distinct classes of non-Abelian braiding statistics have been only studied separately, namely, respectively for MZMs without the need of symmetry protection, whose braiding can be assisted by quantum-dot (QD) Flensberg (2011); Liu et al.
 (2021); Gong et al.
@@
 (2022); Xu et al.
+(2023), the symmetry protected Majorana pairs, and Dirac fermion zero modes.
+Nevertheless, whether the three classes of braiding statistics can be unified in a single quantum model was not known.
+Such a unified model not only is fundamentally significant, but also may show insights into the ultimate experimental realization of the non-Abelian statistics.
+In this article, we show that the minimal Kitaev chain Leijnse and Flensberg (2012); Tsintzis et al.
+(2024) can play the role of such a unified model via constructing proper Y-junction configuration.
+In particular, when the electron hopping amplitude coincides with the superconducting pairing, the minimal Kitaev chain based Y-junction captures the non-Abelian braiding of MZMs (including the QD-assisted Majorana braiding).
+Conversely, when the electron hopping and superconducting pairing are different, the model can realize the non-Abelian braiding of Majorana pairs.
+Notably, when the superconducting pairing amplitude continuously decreases to zero, the non-Abelian Majorana pairs undergo a continuous crossover to non-Abelian Dirac fermion zero modes.
+This result shows that with this unified model the non-Abelian braiding can be realized over a broad experimental parameter regime without fine tuning.
+Particularly, the non-Abelian braiding statistics of the Dirac zero modes can be achieved by relaxing the constraint on fine tuning of the electron hopping and superconducting pairing in the minimal Kitaev chain based Y-junction.
+Our finding highlights that non-trivial physics can also emerge within this newly extended parameter regime, which is of high experimental feasibility and can be implemented using ultracold atoms Jenkins et al.
+(2022) or QD arrays Dvir et al.
+(2023); Bordin et al.
+(2023); Zatelli et al.
+(2024); Bordin et al.
+(2024); ten Haaf et al.
+(2024); Bordin et al.
 (2025).
-In this braiding scheme, the low-lying states involved include both the MZMs and Dirac fermion modes.
-The presence of such braiding scheme intermediating between the Majorana and the Dirac fermionic systems further hints us that these braiding schemes mentioned above might be described in a unified model.
-In this Letter, we demonstrate that the minimal Kitaev chain model Leijnse and Flensberg (2012); Tsintzis et al.
-(2024) can play this role.
-To be specific, under different parameter choices, the minimal Kitaev chain can depict the non-Abelian braiding of MZMs, Dirac fermion zero modes, as well as the QD-assisted Majorana braiding.
-Significantly, the quasiparticles being braided in our setting can switch from the MZMs to Dirac fermion zero modes with a small change in parameters.
-It indicates that the search for non-Abelian anyons shall not be confined to the realm of MZMs.
-In fact, the degenerate Dirac fermion zero modes can always be regarded as the symmetry protected “doubling” of the MZMs Hong et al.
-(2022, 2024).
-In addition to providing new insights into non-Abelian anyons, our proposed study based on minimal Kitaev chain model is experimentally feasible and could be detected using ultracold atoms or QD arrays.
+The latter (QD array) is closely related to the very recent experimental advancements Bordin et al.
+(2024); ten Haaf et al.
+(2024); Bordin et al.
+(2025).
 […]
@@
 (2011); von Oppen et al.
-(2017) jointed by three minimal Kitaev chains [see Fig.
+(2017); Trif and Simon (2019) jointed by three minimal Kitaev chains [see Fig.
 1(a)]
@@
 In general, $\chi$’s and $\Omega$’s are complex as $\chi_{j}=|\chi_{j}|e^{i\theta_{\chi_{j}}}$ and $\Omega_{j}=|\Omega_{j}|e^{i\theta_{\Omega_{j}}}$.
-Significantly, it is impossible to gauge away all these six phases $\theta_{\chi_{j}}$ and $\theta_{\Omega_{j}}$ ($j=1,2,3$), since in Eq.
-(1), there are only four electron operators can be utilized to absorb the gauge phases.
-In other words, the “sweet spot” condition cannot always be satisfied when we consider the braiding in this minimal-Kitaev-chain-based Y-junction.
+Significantly, the “sweet spot” condition is satisfied when $|\chi_{j}|=|\Omega_{j}|$, and $\theta_{\chi_{j}},\theta_{\Omega_{j}}=0$, which requires fining tuning of the experimental parameters.
 Nonetheless, the non-Abelian braiding physics shall not collapse due to a slight deviation of the parameters from the “sweet spot”.
@@
 It implies that this Y-junction can depict the non-Abelian braiding of different kinds of low-lying modes for different choices of $\chi$’s and $\Omega$’s.
-Specifically, there are four distinct cases based on the magnitudes and the phases of $\chi_{j}$ and $\Omega_{j}$, which are summarized in TABLE 1 and will be discussed below in detail.
+Specifically, there are four distinct cases based on the magnitudes and the phases of $\chi_{j}$ and $\Omega_{j}$, which are summarized in TABLE 1.
+The fact that the proposed Y-junction can realize the non-Abelian braiding of Majorana pairs and Dirac fermion zero modes with control, in addition to the single MZMs shall significantly mitigate experimental challenges in realizing the non-Abelian statistics, as studied below in detail.
 $\theta_{\chi_{j}},\theta_{\Omega_{j}}=0$ $\theta_{\chi_{j}},\theta_{\Omega_{j}}\neq 0$
 $|\chi_{j}|=|\Omega_{j}|$ Case I: poorman’s MZMs (PMMs) Case II: QD-assisted MZM braiding
-$|\chi_{j}|\neq|\Omega_{j}|$ Case III: Dirac fermion zero modes (unitary-symmetry preserved) Case IV: Dirac fermion zero modes (unitary-symmetry broken)
+$|\chi_{j}|\neq|\Omega_{j}|$ Case III: unitary-symmetry protected Majorana pairs/Dirac fermion zero modes Case IV: unitary-symmetry broken Majorana pairs/Dirac fermion zero modes
 Table 1: The minimal-Kitaev-chain-based Y-junction can depict the non-Abelian braiding of different types of low-lying modes depending on the choices of $\chi$’s and $\Omega$’s.
+Figure 2: (a) Schematic representation of the Majorana Hamiltonian [Eq.
+(2)] under the conditions that $|\chi_{j}|=|\Omega_{j}|$.
+The Fock space we consider is spanned by six MZMs inside the dashed box.
+(b), (c) The energy spectra during the braiding as (b) Case I in TABLE 1 that $|\chi_{j}|=|\Omega_{j}|$ and $\varphi_{2},\varphi_{3}=0$.
+All the states in (b) are doubly degenerate; (c) Case II in TABLE 1 that $|\chi_{j}|=|\Omega_{j}|$ and $\varphi_{2},\varphi_{3}\neq 0$ ($\varphi_{2}=\pi/6$, and $\varphi_{3}=\pi/12$ here for illustration).
+$|\chi|\equiv\sqrt{|\chi_{1}|^{2}+|\chi_{2}|^{2}+|\chi_{3}|^{2}}$ is chosen as the energy unit.
 The braiding operation in this Y-junction is conducted through coupling the three outer sites with the central site in turn Sau et al.
@@
 In this way, the zero-energy mode initially localized at site 2 is adiabatically moved to site 1.
-In the second step, in the same manner, $\chi_{3}$ and $\Omega_{3}$ are both increased to finite values, and $\chi_{2}$ and $\Omega_{2}$ are decreased to zero in an adiabatic way so that the zero-energy mode in site 3 is moved to site 2.
+Similarly, in the second step, $\chi_{3}$ and $\Omega_{3}$ are both increased to finite values, and $\chi_{2}$ and $\Omega_{2}$ are decreased to zero in an adiabatic way so that the zero-energy mode in site 3 is moved to site 2.
 In the final step, the braiding is accomplished by increasing $\chi_{1}$ and $\Omega_{1}$ to their initial values while decreasing $\chi_{3}$ and $\Omega_{3}$ to zero.
@@
 case I and case II in TABLE 1.
-By decomposing the electron operators $d_{j}$ into Majorana operators $\gamma_{j}$ and $\widetilde{\gamma}_{j}$ ($j=0,1,2,3$), the original Hamiltonian [Eq.
+By decomposing the electron operators $d_{j}$ into Majorana ones $\gamma_{j}$ and $\widetilde{\gamma}_{j}$, the original Hamiltonian [Eq.
 (1)] can be rewritten as Sup :
@@
 $-i|\chi_{3}|(-\sin\varphi_{3}\widetilde{\gamma}_{0}+\cos\varphi_{3}\gamma_{0})\gamma_{3}.$
-Here, as we have stated before, $\varphi_{2}\equiv(\theta_{\chi_{2}}+\theta_{\Omega_{2}}-\theta_{\chi_{1}}-\theta_{\Omega_{1}})/2$ and $\varphi_{3}\equiv(\theta_{\chi_{3}}+\theta_{\Omega_{3}}-\theta_{\chi_{1}}-\theta_{\Omega_{1}})/2$ are two phase constants ($\theta_{\chi_{j}}$ and $\theta_{\Omega_{j}}$ remain unchanged during the braiding) that cannot be absorbed even after gauge transformation Sup .
+Here, $\varphi_{2}\equiv(\theta_{\chi_{2}}+\theta_{\Omega_{2}}-\theta_{\chi_{1}}-\theta_{\Omega_{1}})/2$ and $\varphi_{3}\equiv(\theta_{\chi_{3}}+\theta_{\Omega_{3}}-\theta_{\chi_{1}}-\theta_{\Omega_{1}})/2$ are two phase constants ($\theta_{\chi_{j}}$ and $\theta_{\Omega_{j}}$ remain unchanged during the braiding) that cannot be absorbed even after gauge transformation Sup .
 This Majorana Hamiltonian [Eq.
@@
 2(a)], because fermion parity is conserved here and the Hamiltonian in the parity-even subspace is identical to that in the parity-odd subspace (up to a constant) Sup .
-It is worth noting that though $\widetilde{\gamma}_{1}$ is isolated from the other MZMs, it is still intentionally considered to ensure that the dimension of the Fock space is an integer.
+Although $\widetilde{\gamma}_{1}$ is isolated from the other MZMs, it is still intentionally considered to ensure that the Fock space dimension is an integer.
 In the case I of TABLE 1 that $\varphi_{2},\varphi_{3}=0$, the two-fold degenerate negative-energy states are gapped from the two-fold degenerate positive-energy states (the gap is denoted by $\Delta$), and these degeneracies are preserved throughout the braiding [see Fig.
@@
 The Berry phase during the braiding can be evaluated analytically Sup , which leads to an evolution operator as $U_{\mathrm{case~{}I}}=\begin{pmatrix}e^{-i\pi/4}&0\\ 0&e^{i\pi/4}\end{pmatrix}$.
-This is precisely the Majorana braiding operator Ivanov (2001); von Oppen et al.
-(2017) as expected, since the “sweet spot” condition is satisfied in this case, ensuring that all the low-lying modes in the system are PMMs.
-Figure 2: (a) Schematic representation of the Majorana Hamiltonian [Eq.
-(2)] under the conditions that $|\chi_{j}|=|\Omega_{j}|$.
-The Fock space we consider is spanned by six MZMs inside the dashed box.
-(b), (c) The energy spectra during the braiding as (b) Case I in TABLE 1 that $|\chi_{j}|=|\Omega_{j}|$ and $\varphi_{2},\varphi_{3}=0$.
-All the states in (b) are doubly degenerate; (c) Case II in TABLE 1 that $|\chi_{j}|=|\Omega_{j}|$ and $\varphi_{2},\varphi_{3}\neq 0$ ($\varphi_{2}=\pi/6$, and $\varphi_{3}=\pi/12$ here for illustration).
-$|\chi|\equiv\sqrt{|\chi_{1}|^{2}+|\chi_{2}|^{2}+|\chi_{3}|^{2}}$ is chosen as the energy unit.
+This is precisely the Majorana braiding operator Ivanov (2001) as expected, since the “sweet spot” condition is satisfied in this case, ensuring that all the low-lying modes here are PMMs.
 In the case II of TABLE 1 that $\varphi_{2},\varphi_{3}\neq 0$, the two-fold ground-state degeneracy is lifted except for three special parameter points during the braiding [see Fig.
 2(c)].
-Therefore, in addition to the Berry phase, the dynamic phase part in the evolution operator $U=\hat{P}\exp[\int-iE(t)\mathrm{d}t+i\mathbf{A}(\boldsymbol{\lambda})\cdot\mathrm{d}\boldsymbol{\lambda}]$ Wilczek and Zee (1984); Chang (2021) generally cannot be dropped.
+Therefore, in addition to the Berry phase, the dynamic phase in the evolution operator $U=\hat{P}\exp[\int-iE(t)\mathrm{d}t+i\mathbf{A}(\boldsymbol{\lambda})\cdot\mathrm{d}\boldsymbol{\lambda}]$ Wilczek and Zee (1984); Chang (2021) generally cannot be dropped.
 Here, $\hat{P}$ represents the time-ordering operator, $E$ is a diagonal matrix corresponding to the eigenenergy spectra, and $\mathbf{A}$ is the Berry connection.
-Under the adiabatic condition, where $T\gg 1/\Delta$, we could still focus only on the two lowest states with negative energies [see Fig.
+Under the adiabatic condition $T\gg 1/\Delta$, we could still focus only on the two lowest states with negative energies [see Fig.
 2(c)].
@@
 Therefore, the minimal Kitaev chain here exactly describes the QD-assisted braiding that two MZMs are swapped via a Dirac fermion zero mode.
-Non-Abelian braiding of Dirac fermion zero modes.
-The more intriguing scenario is that even though the equalities $|\chi_{j}|=|\Omega_{j}|$ ($j=1,2,3$) are only slightly violated, the low-lying modes being braided are no longer single MZMs, but Majorana pairs.
+Non-Abelian braiding of Majorana pairs and Dirac fermion zero modes.
+The more intriguing scenario is that even though the equalities $|\chi_{j}|=|\Omega_{j}|$ ($j=1,2,3$) are only slightly violated (an unavoidable circumstance in realistic condition), the low-lying modes being braided are no longer single MZMs, but Majorana pairs.
 To be specific, under the conditions $|\chi_{j}|\neq|\Omega_{j}|$ ($j=1,2,3$), i.e.
@@
 Therefore, the low-lying modes being braided here are actually fully-overlapped Majorana pairs respecting local unitary symmetry $\mathcal{R}$.
-Then each Majorana pair as a whole is actually Dirac fermion zero mode Hong et al.
-(2022); Wu et al.
-(2020a, 2023) that the corresponding braiding properties can be constructed from the Majorana ones.
-In the adiabatic limit $T\gg\frac{1}{w},\frac{1}{\Delta}$ [$w$ is the (average) energy difference between the lowest two-fold states and the second lowest two-fold states, see Fig.
+Then each Majorana pair as a whole is actually complex fermion mode Wu et al.
+(2023) that the corresponding braiding properties can be constructed from the Majorana ones.
+In the adiabatic limit $T\gg 1/w,1/\Delta$ [$w$ is the (average) energy difference between the lowest two-fold states and the second lowest two-fold states, see Fig.
 3(c)], the Berry phase for the lowest two-fold states turns out to be the projection of the tensor product of the Majorana one’s into the parity-even subspace as ^$\hat{\mathbb{P}}_{\mathrm{even}}[\mathrm{diag}\{e^{-i\pi/4},e^{i\pi/4}\}\otimes\mathrm{diag}\{e^{-i\pi/4},e^{i\pi/4}\}]=\mathrm{diag}\{-i,i\}$.
 Similarly, the Berry phase for the second lowest two-fold states is ^$\hat{\mathbb{P}}_{\mathrm{even}}[\mathrm{diag}\{e^{-i\pi/4},e^{i\pi/4}\}\otimes\mathrm{diag}\{e^{i\pi/4},e^{-i\pi/4}\}]=\mathrm{diag}\{1,1\}$ Sup .
-In addition to this, the dynamic phases also contribute when $T\gg\frac{1}{w},\frac{1}{\Delta}$, and hence the braiding operator reads $U_{\mathrm{case~{}III}}=e^{i\theta}\left[e^{-i\delta}\begin{pmatrix}-i&0\\ 0&i\end{pmatrix}\oplus e^{i\delta}\begin{pmatrix}1&0\\ 0&1\end{pmatrix}\right]$.
-Here, $\theta$ ($\delta$) are the dynamic phases corresponding to the average energy of (energy difference between) the lowest and second lowest two-fold dengenerate states.
-By contrast, for $1/w\gg T\gg 1/\Delta$, the dynamic phase of the four lowest states only contributes a trivial overall phase and can be dropped.
-Moreover, the Berry phase for the Majorana set $\{\widetilde{\gamma}_{j}\}$ now reduces to an identity $I_{4\times 4}$ since $\widetilde{\gamma}_{j}$’s are braided in the non-adiabatic limit $1/w\gg T$ Chang (2021).
-Therefore, in this case, the braiding operator is obtained by projecting the tensor product $\mathrm{diag}\{e^{-i\pi/4},e^{i\pi/4}\}\otimes I_{4\times 4}$ into the parity-even subspace, giving $U_{\mathrm{case~{}III}}=\mathrm{diag}\{e^{-i\pi/4},e^{-i\pi/4},e^{i\pi/4},e^{i\pi/4}\}$ Sup .
+Meanwhile, the dynamic phases also contribute.
+Hence, the braiding operator reads $U_{\mathrm{case~{}III}}=e^{i\theta}\left[e^{-i\delta}\begin{pmatrix}-i&0\\ 0&i\end{pmatrix}\oplus e^{i\delta}\begin{pmatrix}1&0\\ 0&1\end{pmatrix}\right]$, where $\theta$ ($\delta$) are the dynamic phases corresponding to the average energy of (energy difference between) the lowest and second lowest two-fold degenerate states.
+By contrast, for $1/w\gg T\gg 1/\Delta$, the dynamic phase of the four lowest states only contributes a trivial overall phase.
+Moreover, the Berry phase for the Majorana set $\{\widetilde{\gamma}_{j}\}$ is an identity $I_{4\times 4}$ here since $\widetilde{\gamma}_{j}$’s are braided in the non-adiabatic limit $1/w\gg T$ Chang (2021).
+Therefore, in this case, the braiding operator is obtained by projecting $\mathrm{diag}\{e^{-i\pi/4},e^{i\pi/4}\}\otimes I_{4\times 4}$ into the parity-even subspace, giving $U_{\mathrm{case~{}III}}=\mathrm{diag}\{e^{-i\pi/4},e^{-i\pi/4},e^{i\pi/4},e^{i\pi/4}\}$ Sup .
 Figure 3: (a), (b) Schematic representation of the Majorana Hamiltonian [Eq.
@@
 This is in analogy with the consequence stated above that the Berry phase is the same for case II and case I.
-The difference lies in that now the dynamic phases $\theta\mp\delta$ are $\varphi_{2}$-, $\varphi_{3}$-, and $|\Omega_{j}|/|\chi_{j}|$-dependent since the energy during the braiding vary with $\varphi_{2}$, $\varphi_{3}$, and $|\Omega_{j}|/|\chi_{j}|$.
+The difference lies in that now the dynamic phases $\theta\mp\delta$ are $\varphi_{2}$-, $\varphi_{3}$-, and $|\Omega_{j}|/|\chi_{j}|$-dependent since the energy spectra vary with $\varphi_{2}$, $\varphi_{3}$, and $|\Omega_{j}|/|\chi_{j}|$.
 Meanwhile, the braiding results still exhibit an oscillation with increasing $T$ Sup .
@@
 In this way, the $1/w\gg T\gg 1/\Delta$ condition can no longer be satisfied when $\varphi_{2}$, $\varphi_{3}$, or $\big{|}|\chi_{j}|-|\Omega_{j}|\big{|}$ is significantly non-zero.
-As a result, an additional dynamic phase is always accumulated alongside the geometric phase, providing a clear signature of the symmetry-breaking effect Wu et al.
-(2020a, 2023, b) on the braiding of Dirac fermion zero modes.
+Consequently, an additional dynamic phase is always accumulated alongside the geometric phase, a clear signature of the symmetry-breaking effect Wu et al.
+(2020b) on the braiding of Dirac fermion zero modes.
 Discussions.
-The minimal Kitaev chain model is experimentally feasible, in ultracold atom systems for instance, especially for $|\Omega_{j}|=0\neq|\chi_{j}|$, as the complexity of pairing term is circumvented.
-In this case, the low-lying modes are Dirac fermion modes with unitary symmetry preserved (case III) or broken (case IV), other than MZMs.
-In this condition, the Hamiltonian in Eq.
-(1) can be engineered by utilizing the internal degree of freedom of atoms.
+The “sweet spot” condition has long been a key objective in experiments Bordin et al.
+(2024); ten Haaf et al.
+(2024); Bordin et al.
+(2025); Liu et al.
+(2022).
+We have shown here the non-trivial phases with symmetry protected MZM pairs or Dirac zero modes in the regime away from this condition.
+With the removal of fine-tuning condition $|\Omega_{j}|=|\chi_{j}|$, the minimal Kitaev chain model becomes more realistic for experimental realization.
+Particularly, our prediction is of high experimental feasibility by setting $|\chi_{j}|>|\Omega_{j}|=0$, as the complexity associated with the superconducting pairing is circumvented.
+In this case, the low-lying modes being braided are Dirac fermion modes that the fermion number is conserved during the braiding.
+The minimal Kitaev chain has recently been experimentally engineered in QD arrays Bordin et al.
+(2024); ten Haaf et al.
+(2024); Bordin et al.
+(2025) and could also be implemented by utilizing the internal degrees of freedom of ultracold atoms.
 Taking 171Yb fermions as an example, the four braiding modes correspond to the atomic spin states, 1S0$|F=1/2,m_{F}=\pm 1/2\rangle$, and ${}^{3}P$0$|F=1/2,m_{F}=\pm 1/2\rangle$, which have significantly long lifetime.
-The tunnelings among the four fermion modes are highly controllable by manipulating Raman and clock transitions Jenkins et al.
+The tunnelings among the four fermion modes are highly controllable and readily realized by manipulating Raman and clock transitions Jenkins et al.
 (2022).
+This study paves the way for the ultimate realization of non-Abelian statistics in experiment.
 Acknowledgements.
+The authors thank Leo P.
+Kouwenhoven and Chun-Xiao Liu for fruitful discussions.
 This work was financially supported by the Innovation Program for Quantum Science and Technology (Grant No.
@@
 12425401 and No.
-12261160368), National Key Research and Development Program of China (2021YFA1400900), and by Shanghai Municipal Science and Technology (Grant No.
-24DP2600100).
+12261160368), National Key Research and Development Program of China (Grant No.
+2021YFA1400900 and No.
+2024YFA1409000), Shanghai Municipal Science and Technology (Grant No.
+24DP2600100), Shanghai Municipal Science and Technology Major Project (Grant No.2019SHZDZX01), and by Shanghai Science and Technology Innovation Action Plan (Grant No.
+24LZ1400800).